# Harvesting primordial black holes from stochastic trees with FOREST

### v1

@@ 1 Introduction
This framework has been used to study the statistics of large fluctuations, which have been shown to feature exponential tails [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. These are substantially heavier than Gaussian tails, which leads to more abundant large-curvature regions than what cosmological perturbation theory would predict. However, determining the real-space profile of the curvature perturbation within the stochastic formalism remains challenging. This difficulty arises because, when simulating realizations of the Langevin equation that governs the stochastic dynamics of the background fields, only a single worldline is tracked until inflation ceases. Repeating this procedure generates an ensemble of such worldlines, but it provides no information about the spatial arrangement of the Hubble patches that terminate them on the end-of-inflation hypersurface.
@@
Note that other numerical approaches can be followed to simulate non-perturbative dynamics during inflation, such as numerical relativistic codes [76, 77, 78, 79, 80] (mostly employed to investigate the onset of inflation from non-homogeneous initial conditions) or lattice codes [81, 82, 83, 84, 85]. These approaches are numerically expensive, and do not include the effect of quantum diffusion, which is nonetheless essential in predicting the statistics of large fluctuations during inflation. Recently however, lattice codes have been extended to include stochastic noise and perform the stochastic-$\delta N$ program. In particular, the code STOLAS [86] delivers three-dimensional maps of the curvature perturbation, where the effect of quantum diffusion and of non-linear evolution is included at large scales. It thus represents a significant advance.
@@ 2 Stochastic trees for inflation
In this expression, $V(\phi)$ is the potential energy stored in the inflaton, and $\xi(N)$ is a normalized white Gaussian noise, such that $\expectationvalue{\xi(N)}=0$ and $\expectationvalue{\xi(N)\xi(N^{\prime})}=\delta(N-N^{\prime})$. Along the slow-roll attractor the effective phase-space is one-dimensional [96], but the extension of stochastic trees to higher-dimensional systems is straightforward.
@@ 2.1 The tree structure of inflating spacetimes
If inflation never stops, the tree grows indefinitely. In practice however, inflation has to stop and the nodes where this occurs correspond to the “leaves” of the tree. In the tree sketched in fig. 2, the leaves are the nodes with labels $4$, $20$, $21$, $11$, $6$, $14$ and $15$. In single-field slow-roll models, inflation most often stops by slow-roll violation, i.e. when $\phi$ reaches a certain field value $\phi_{\mathrm{end}}$ where the potential function becomes too steep for inflation to proceed. This is why, when evolving $\phi$ from a parent node $i$ to one of its child nodes $\ell$, if $\phi$ crosses $\phi_{\mathrm{end}}$ then the process is terminated and a leaf $L_{\ell}$ is created. In units of the Hubble volume $V_{\sigma}=4\pi R_{\sigma}^{3}/3$ at the end of inflation, the physical volume of that leaf is given by
$V\left(L_{\ell}\right)=e^{3\left(\mathcal{N}_{i\to\ell}-\Delta N\right)}\in[1,2[\,,$ (2.3)
@@
Stochastic trees of that type can be efficiently generated numerically using recursive methods, since each vertex gives rise to two subtrees that can be processed independently. Once the routine implementing the elementary vertex depicted in fig. 1 is programmed, it shall call itself from each child node, until inflation ends at the tip of every branch and the entire tree is scanned. We have developed FOREST (FOrtran Recursive Exploration of Stochastic Trees), a parallel Modern Fortran code using MPI+OpenMP that proceeds along these lines, and which is used to produce all numerical results presented in this paper. FOREST takes as an input an inflationary potential function $V(\phi)$, its derivative $V^{\prime}(\phi)$, and boundary conditions for the field configuration. The Langevin equation (2.1) is solved using the Euler–Maruyama method [97] with a varying step $\delta N$. Close to the end-of-inflation absorbing boundary, we enforce $\delta N<3[2\pi(\phi-\phi_{\mathrm{end}})]^{2}/V(\phi)/5$ to limit the probability of barrier crossing to $5\sigma$ and avoid double crossings that spoil estimations of the first-passage time (see appendix B for more details). The value of the Hubble function is updated at each step $\delta N$ using Friedmann’s equations. We have successfully tested FOREST with populations of more than $10^{12}$ stochastic trees, shared across $2\times 256$ cores using MPI+OpenMP. We report excellent scaling as a function of the number of cores as well as the number of trees (see appendix A). We intend to make FOREST publicly accessible soon after publication.
@@ 2.3 Mapping the curvature perturbation in comoving space
Figure 4: Example of a two-dimensional comoving map with five PBHs in the flat-well toy model discussed in section 4, with $\mu=0.8$ and starting from $\phi=\Delta\phi_{\mathrm{well}}$. The colour scale indicates the elapsed number of $e$-folds of the different patches on the end-of-inflation hypersurface. This is the same realization as in fig. 5. On the right panel, we mark the five PBHs by superimposing black patches, according to the PBH formation criterion presented in section 3.2. We single out this realization for illustrative and pedagogical purposes, it is not particularly representative of the tree population, in which PBH formation remains a rare event.
@@ 3.1 Coarse-shelled curvature perturbation
This “coarse-shelled” curvature perturbation can be related to the Laplacian of the curvature perturbation, or more generally to the compaction function, as follows. The region emerging from node $i$ is composed of the regions stemming from its two child nodes $\ell$ and $m$, but the geometry of these subregions is not specified. 222 When constructing the maps of section 2.3, a choice was made that consists in splitting volumes along orthogonal planes, the orientation of which alternates as the time elapsed from the primeval patch increases, such that statistical isotropy is recovered at large scales. That choice is however arbitrary, it is a discretization convention. Only when averaged over large volumes can discretization artefacts disappear, see section 5.3. Since the Laplacian is a spherically symmetric operator, let us picture the regions emerging from nodes $i$ and $\ell$ as being two concentric spheres. In this case, $\zeta_{\ell i}$ is the curvature perturbation coarse grained over the shell located between the spheres of radius $R_{\ell}$ and $R_{i}$, where $V_{i}=4\pi R_{i}^{3}/3$ and likewise for $V_{\ell}$, which is why it is called the “coarse-shelled” curvature perturbation. For the noise appearing in the Langevin equation (2.1) to be white, i.e. uncorrelated over time, coarse graining has to be performed by means of a sharp window function in Fourier space, $\zeta_{R}(\vec{k})=\theta(a/R-k)\zeta(\vec{k})$, where $\theta(x)$ is the Heaviside function. In this case,
@@ 3.2 Unbalance index
Figure 5: Example of a tree realization with five PBHs in the flat-well toy model, with $\mu=0.8$ and starting from $\phi=\Delta\phi_{\mathrm{well}}$. This is the same realization as fig. 4. The black patches are those for which the coarse-shelled curvature perturbation, used as a proxy for the compaction function, exceeds the PBH formation threshold.
@@ 5.3 Discretization artefacts
For the sake of simplicity, we focus on a light test scalar field in a fixed binary tree. [^6] Instead of a binary tree, one may consider trees with more branches attached to each node (for instance, trees where Hubble patches split into $8$ patches after $\Delta N=\ln(2)$ $e$-folds in $3$ dimensions). This would not not change the nature of the arguments developed here. The field being test, inflation ends at the same time in all the branches, i.e. the tree is perfectly balanced. This makes it easier to relate topological distances in the tree and physical distances on the final hypersurface. Although it may be seen as restrictive, it will allow us to show explicitly how discretization artefacts disappear at large scales, and to argue why this is in fact a generic result.
@@ 5.3.1 Branching times
In stochastic trees, branching is an instantaneous event, but the precise time at which it takes place is not unambiguous. For instance, instead of splitting a node $i$ and letting the branches leading to the nodes $\ell$ and $m$ grow independently, one could let the node $i$ grow to twice its volume and then split it. More generally, one could consider a hybrid process where the node $i$ grows up to a time $\alpha\Delta N$, then splits, and the two child branches are evolved independently for $(1-\alpha)\Delta N$, where $0\leq\alpha\leq 1$. These three possibilities are depicted in fig. 9. The “split-then-grow” procedure corresponds to $\alpha=0$, while the “grow-then-split” procedure corresponds to $\alpha=1$.
@@
The paths leading to $i$ and $j$ coincide until node $n_{*}$, and then become independent. In the example depicted in fig. 10, for $i=20$ and $j=23$, one has $q_{\mathrm{T}}=4$, $d=3$ and the last common ancestor has label $n_{*}=5$. First, consider the case where, from $n_{*}$, the path leading to $i$ always proceeds leftwards, while the path leading to $j$ always proceeds rightwards (we will show how to go beyond this assumption in section 5.3.2). In this case, the last common ancestor has label $n_{*}=i/(d+1)$.777 More generally, it is convenient to use the binary representation of the nodes to identify the last common ancestor. For example, $i=20=10100_{2}$ and $j=23=10111_{2}$, their last common ancestor is the longest sequence of identical bits in the two numbers, so $5101_{2}=$ 5. The duration of the common path between nodes $i$ and $j$ is thus
@@ 5.3.2 Branching surfaces
This shows that, even though $\Sigma(d_{\mathrm{P}})$ receives contribution from pairs of leaves with $\ln(\sigma Hd_{\mathrm{P}})\leq\Delta N_{*}(i,j)\leq N_{\mathrm{T}}$, in practice those with $\Delta N_{*}(i,j)\simeq\ln(\sigma Hd_{\mathrm{P}})$ dominate, and the correct result is recovered at large distance. The discretization scheme could be improved, for instance by randomly drawing the position of the branching hypersurfaces within the patches. This was proposed in Ref. [70] and makes the trees statistically homogeneous indeed. However, it never produces more correctly correlated pairs than the standard tree, hence it does not improve the convergence towards the correct result at large distance.
@@ 6 Conclusion
This intricate structure can be numerically realized with FOREST, a cutting-edge, parallel code that generates and scans vast populations of stochastic trees to deliver relevant statistics. We showed how FOREST returns the first-passage time, in terms of $e$-folds, through the end-of-inflation hypersurface, averaged over the volume emerging from the initial node, and how the coarse-grained curvature perturbation at the Hubble scale at the end of inflation, or at any arbitrary scale, can be reconstructed.
@@
By simulating stochastic trees, FOREST reconstructs statistics through direct sampling, enabling, for instance, the investigation of the first-passage-time distribution down to values of probability as small as $10^{-10}$, representing a significant improvement over statistics inferred from simple Langevin simulations and lattice codes. To investigate the tails of the distributions even more efficiently, methods based on importance sampling can be employed [116, 95], and we leave this possibility for future versions of FOREST.
@@ Acknowledgments
The authors thank Marie-Paule Faber, Christophe Ringeval and Juan García Bellido for insightful discussions. Pierre Auclair is a Postdoctoral Researcher of the Fonds de la Recherche Scientifique – FNRS. Baptiste Blachier is a Research Fellow of the Fonds de la Recherche Scientifique – FNRS. Chiara Animali is supported by the ESA Belgian Federal PRODEX Grant $\mathrm{N^{\circ}}4000143201$. Computational resources have been provided by the CURL’s development cluster, the supercomputing facilities of the Université catholique de Louvain (CISM/UCL) and the Consortium des Équipements de Calcul Intensif en Fédération Wallonie Bruxelles (CÉCI) funded by the Fond de la Recherche Scientifique de Belgique (F.R.S.-FNRS) under convention 2.5020.11 and by the Walloon Region.

### v2

@@ 1 Introduction
This framework has been used to study the statistics of large fluctuations, which have been shown to feature exponential tails [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. These are substantially heavier than Gaussian tails, which leads to more abundant large-curvature regions than what cosmological perturbation theory would predict. However, determining the real-space profile of the curvature perturbation within the stochastic formalism remains challenging. This difficulty arises because, when simulating realizations of the Langevin equation that governs the stochastic dynamics of the background fields, only a single worldline is tracked until inflation ceases. Repeating this procedure generates an ensemble of such worldlines, but it provides no information about the spatial arrangement of the Hubble patches that terminate them on the end-of-inflation hypersurface.
@@
Note that other numerical approaches can be followed to simulate non-perturbative dynamics during inflation, such as numerical relativistic methods [77, 78, 79, 80, 81, 82] (mostly employed to investigate the onset of inflation from non-homogeneous initial conditions) or lattice codes [83, 84, 85, 86, 87].
These approaches are numerically expensive, and do not include the effect of quantum diffusion, which is nonetheless essential in predicting the statistics of large fluctuations during inflation. Recently however, lattice codes have been extended to include stochastic noise and perform the stochastic-$\delta N$ program. In particular, the code STOLAS [88] delivers three-dimensional maps of the curvature perturbation, where the effect of quantum diffusion and of non-linear evolution is included at large scales. It thus represents a significant advance.
@@ 2 Stochastic trees for inflation
In this expression, where the number of $e$-folds $N$ is used as a time variable [98], $V(\phi)$ is the potential energy stored in the inflaton, and $\xi(N)$ is a normalized white Gaussian noise, such that $\expectationvalue{\xi(N)}=0$ and $\expectationvalue{\xi(N)\xi(N^{\prime})}=\delta(N-N^{\prime})$. Along the slow-roll attractor the effective phase-space is one-dimensional [99], but the extension of stochastic trees to higher-dimensional systems is straightforward.
@@ 2.1 The tree structure of inflating spacetimes
If inflation never stops, the tree grows indefinitely. In practice however, inflation has to stop and the nodes where this occurs correspond to the “leaves” of the tree. In the tree sketched in fig. 2, the leaves are the nodes with labels $4$, $20$, $21$, $11$, $6$, $14$ and $15$. In single-field slow-roll models, inflation most often stops by slow-roll violation, i.e. when $\phi$ reaches a certain field value $\phi_{\mathrm{end}}$ where the potential function becomes too steep for inflation to proceed. This is why, when evolving $\phi$ from a parent node $i$ to one of its child nodes $\ell$, if $\phi$ crosses $\phi_{\mathrm{end}}$ then the process is terminated and a leaf $L_{\ell}$ is created. In units of the Hubble volume $V_{\sigma}=4\pi R_{\sigma}^{3}/3$ at the end of inflation, the physical volume of that leaf is given by (see footnote 7)
$V\left(L_{\ell}\right)=e^{3\left(\mathcal{N}_{i\to\ell}-\Delta N\right)}\in[1/2,1[\,,$ (2.3)
@@
Stochastic trees of that type can be efficiently generated numerically using recursive methods, since each vertex gives rise to two subtrees that can be processed independently. Once the routine implementing the elementary vertex depicted in fig. 1 is programmed, it shall call itself from each child node, until inflation ends at the tip of every branch and the entire tree is scanned. We have developed FOREST (FOrtran Recursive Exploration of Stochastic Trees), a parallel Modern Fortran code using MPI+OpenMP that proceeds along these lines, and which is used to produce all numerical results presented in this paper. FOREST takes as an input an inflationary potential function $V(\phi)$, its derivative $V^{\prime}(\phi)$, and boundary conditions for the field configuration. The Langevin equation (2.1) is solved using the Euler–Maruyama method [100] with a varying step $\delta N$. Close to the end-of-inflation absorbing boundary, we enforce $\delta N<3[2\pi(\phi-\phi_{\mathrm{end}})]^{2}/V(\phi)/5$ to limit the probability of barrier crossing to $5\sigma$ and avoid double crossings that spoil estimations of the first-passage time (see appendix B for more details). The value of the Hubble function is updated at each step $\delta N$ using Friedmann’s equations. We have successfully tested FOREST with populations of more than $10^{12}$ stochastic trees, shared across $2\times 256$ cores using MPI+OpenMP. We report excellent scaling as a function of the number of cores as well as the number of trees (see appendix A). FOREST is publicly accessible under GPLv3 license [101].
@@ 2.3 Mapping the curvature perturbation in comoving space
Figure 4: Example of a two-dimensional comoving map with five PBHs in the flat-well toy model discussed in section 4, with $\mu=0.8$ and starting from $\phi=\Delta\phi$. The colour scale indicates the elapsed number of $e$-folds of the different patches on the end-of-inflation hypersurface. This is the same realization as in fig. 5. On the right panel, we mark the five PBHs by superimposing black patches, according to the PBH formation criterion presented in section 3.2. We single out this realization for illustrative and pedagogical purposes, it is not particularly representative of the tree population, in which PBH formation remains a rare event.
@@ 3.1 Coarse-shelled curvature perturbation
This “coarse-shelled” curvature perturbation can be related to the Laplacian of the curvature perturbation, or more generally to the compaction function, as follows. The region emerging from node $i$ is composed of the regions stemming from its two child nodes $\ell$ and $m$, but the geometry of these subregions is not specified.222 When constructing the maps of section 2.3, a choice was made that consists in splitting volumes along orthogonal planes, the orientation of which alternates as the time elapsed from the primeval patch increases, such that statistical isotropy is recovered at large scales, see section 5.3. It can thus be used to compute quantities coarse-grained over large volumes. The coarse-shelled curvature perturbation being a local quantity, one should rather employ a scheme that is isotropic already at the level of an individual leave and its parent. Note that if the “concentric-sphere” scheme were embedded globally it would lead to some Hubble patches being pictured as thin layers of large radius, which would be incompatible with the geometry of inflating space-times. Therefore, different schemes should be employed for local and global quantities, such that the space-time symmetries are correctly implemented at both levels. Since the Laplacian is a spherically symmetric operator, let us picture the regions emerging from nodes $i$ and $\ell$ as being two concentric spheres. In this case, $\zeta_{\ell i}$ is the curvature perturbation coarse grained over the shell located between the spheres of radius $R_{\ell}$ and $R_{i}$, where $V_{i}=4\pi R_{i}^{3}/3$ and likewise for $V_{\ell}$, which is why it is called the “coarse-shelled” curvature perturbation. For the noise appearing in the Langevin equation (2.1) to be white, i.e. uncorrelated over time, coarse graining has to be performed by means of a sharp window function in Fourier space, $\zeta_{R}(\vec{k})=\theta(a/R-k)\zeta(\vec{k})$, where $\theta(x)$ is the Heaviside function. In this case,
@@ 3.2 Unbalance index
Figure 5: Example of a tree realization with five PBHs in the flat-well toy model, with $\mu=0.8$ and starting from $\phi=\Delta\phi$. This is the same realization as fig. 4. The black patches are those for which the coarse-shelled curvature perturbation, used as a proxy for the compaction function, exceeds the PBH formation threshold.
@@ 5.3 Discretization artefacts
For the sake of simplicity, we focus on a light test scalar field in a fixed binary tree. [^6] Instead of a binary tree, one may consider trees with more branches attached to each node (for instance, trees where Hubble patches split into $8$ patches after $\Delta N=\ln(2)$ $e$-folds in $3$ dimensions). This would not change the nature of the arguments developed here. The field being test, inflation ends at the same time in all the branches, i.e. the tree is perfectly balanced. This makes it easier to relate topological distances in the tree and physical distances on the final hypersurface. Although it may be seen as restrictive, it will allow us to show explicitly how discretization artefacts disappear at large scales, and to argue why this is in fact a generic result.
@@ 5.3.1 Branching times
[…]
@@
The paths leading to $i$ and $j$ coincide until node $n_{*}$, and then become independent. In the example depicted in fig. 10, for $i=20$ and $j=23$, one has $q_{\mathrm{T}}=4$, $d=3$ and the last common ancestor has label $n_{*}=5$. First, consider the case where, from $n_{*}$, the path leading to $i$ always proceeds leftwards, while the path leading to $j$ always proceeds rightwards (we will show how to go beyond this assumption in section 5.3.2). In this case, the last common ancestor has label $n_{*}=i/(d+1)$.888 More generally, it is convenient to use the binary representation of the nodes to identify the last common ancestor. For example, $i=20=10100_{2}$ and $j=23=10111_{2}$, their last common ancestor is the longest sequence of identical bits in the two numbers, so $5101_{2}=$ 5. The duration of the common path between nodes $i$ and $j$ is thus
@@ 5.3.2 Branching surfaces
This shows that, even though $\Sigma(d_{\mathrm{P}})$ receives contribution from pairs of leaves with $\ln(\sigma Hd_{\mathrm{P}})\leq\Delta N_{*}(i,j)\leq N_{\mathrm{T}}$, in practice those with $\Delta N_{*}(i,j)\simeq\ln(\sigma Hd_{\mathrm{P}})$ dominate, and the correct result is recovered at large distance. The discretization scheme could be improved, for instance by randomly drawing the position of the branching hypersurfaces within the patches. This was proposed in Ref. [71] and makes the trees statistically homogeneous indeed. However, it never produces more “correctly correlated” pairs than the standard tree, hence it does not improve the convergence towards the correct result at large distance.
Note that in this work, we have not computed the two-point correlation functions discussed above, but rather consider one-point distributions of the coarse-grained or coarse-shelled curvature perturbation. Since the method outlined in section 3.1 allows one to perform coarse-shelling between two nodes that are direct siblings only, only such pairs of nodes are accounted for in our stochastic averages. Given that these pairs are “correctly correlated” in the language developed above, we expect our results for the statistics of PBHs to be even more robust to discretisation artefacts than the two-point correlation function. In any case, in the simple example discussed here (a test free field), Gaussian statistics allows one to express the one-point distribution entirely in terms of the two-point correlation functions, hence the convergence of the latter under proper averaging ensures the convergence of the former.
@@ 6 Conclusion
This intricate structure can be numerically realized with FOREST [101], a cutting-edge, parallel code that generates and scans vast populations of stochastic trees to deliver relevant statistics. We showed how FOREST returns the first-passage time, in terms of $e$-folds, through the end-of-inflation hypersurface, averaged over the volume emerging from the initial node, and how the coarse-grained curvature perturbation at the Hubble scale at the end of inflation, or at any arbitrary scale, can be reconstructed.
@@
By simulating stochastic trees, FOREST reconstructs statistics through direct sampling, enabling, for instance, the investigation of the first-passage-time distribution down to values of probability as small as $10^{-10}$, improving on the numerical efficiency of direct Langevin simulations and lattice codes. To investigate the tails of the distributions even more efficiently, methods based on importance sampling can be employed [120, 97], and we leave this possibility for future versions of FOREST.
@@ Acknowledgments
The authors thank Marie-Paule Faber, Christophe Ringeval and Juan García-Bellido for insightful discussions. Pierre Auclair is a Postdoctoral Researcher of the Fonds de la Recherche Scientifique – FNRS. Baptiste Blachier is a Research Fellow of the Fonds de la Recherche Scientifique – FNRS. Chiara Animali is supported by the ESA Belgian Federal PRODEX Grant $\mathrm{N^{\circ}}4000143201$. Computational resources have been provided by the CURL’s development cluster, the supercomputing facilities of the Université catholique de Louvain (CISM/UCL) and the Consortium des Équipements de Calcul Intensif en Fédération Wallonie Bruxelles (CÉCI) funded by the Fond de la Recherche Scientifique de Belgique (F.R.S.-FNRS) under convention 2.5020.11 and by the Walloon Region.